ROTATIONAL SPECTROSCOPY OF SINGLE CARBONYL SULFIDE MOLECULES EMBEDDED IN SU-PERFLUID HELIUM NANODROPLETS

\underline{PAUL L. RASTON}, RUDOLF LEHNIG, AND WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, Alberta T6G-2G2, Canada.

The pure rotation spectrum of carbonyl sulfide embedded in superfluid helium nanodroplets was measured in the frequency range from 4 to 15.5 GHz. Four lines, corresponding to the J = 1-0, J = 2-1, J = 3-2, and J = 4-3 transitions were found. The line widths increase with J, and are about twice that of the corresponding rovibrational lines^{a}, indicating faster rotational relaxation within the ground vibrational manifold. The comparison of the pure rotational spectrum with the microwave-infrared double resonance spectrum^{b} reveals that the double resonance measurement mainly probes rotational transitions within the upper vibrational manifold.

aS. Grebenev, M. Hartmann, M. Havenith, B. Sartakov, J. P. Toennies, and A. F. Vilesov, J. Chem. Phys., 112, 4485 (2000).

bS. Grebenev, M. Havenith, F. Madeja, J. P. Toennies, and A. F. Vilesov, J. Chem. Phys., 113, 9060 (2000).